OpenStudy (anonymous):

sqrt{5x+29=x+3} sqrt{5x+29=x+3} @Mathematics

OpenStudy (anonymous):

\[\sqrt{5x+29=x+3}\]

OpenStudy (anonymous):

Would you like to type it one more time? Probably sqrt only for the left part?

OpenStudy (anonymous):

\[\sqrt{9x+81=x+5}\]

OpenStudy (anonymous):

sorry i had already finished that other one. this is a new one

OpenStudy (anonymous):

It's OK. But you still didn't type the second one correct.

OpenStudy (anonymous):

the line just kept going

OpenStudy (anonymous):

\[\sqrt{9x+81} =x+5\]

jhonyy9 (jhonyy9):

9x+81=x2+10x+25 x2 +x -56 =0 x_1,_2 =(-1+/- sqrt(1+224))/2 =(-1+/- 15)/2 = -8 and 7